Subject: Summer 2000 shutdown

From: Mark Klaene

Submitted: Fri, 28 Jul 2000 09:38:28 -0600

Message number: 447 (previous: 446, next: 448 up: Index)

The following items were accomplished during the recent 3 week summer 
shutdown. While we are still waiting for some clear nights to test 
all of the work that was done this will give everyone a feel for what 
was accomplished by the site staff and folks at UW.

1. M1 was removed, washed, and inspected. The inspection revealed no 
changes from last year.

2. Linear Encoders were install on M3 for the axial supports

3. Pucks were installed on M1 where the axial load cells rest

4. M1 LVDT's were found to be sticking and replaced and calibrated

5. M1 load cells were calibrated

6. A sector load cell on M1 was found loose and repaired

7. B sector load cell on M1  was found bad and replaced

8. The error of the M3 stepper motors not ending on full steps was repaired

9. Repaired the M3 rotational problem with the slot sensor

10. Further test were accomplished towards automating the mirror covers

11. Numerous hard points between the building and the telescope pier 
were replaced with flex

12. The temperature wiring on the telescope truss was replaced with a 
shielded cable

13. The PMSS pumps were serviced

14. The main Az bearing was serviced

15. Changes were made to the M1 ventilation system to eliminate 
forces on the mirror

16.  A new TCC version was installed

17.  Electrical insulators were installed in the cooling lines for Spicam

18.  All Instruments dewars were serviced

19.  Modified dust covers for increased airflow around M3

20.  Collimation and PMSS tuning were improved.

We appreciate the user's patience during this time when the telescope 
was not available but believe that we are returning a telescope to 
you that will provide improved performance in the future. As always 
there may be a few bugs during the initial days following the 
shutdown as once again weather has kept us from fully ringing out the 
telescope prior to returning to science.

Mark Klaene						Po Box 59
Deputy Site Manager					Sunspot NM 88349
Apache Point Observatory				(505) 437-6822

APO APO APO APO APO  Apache Point Observatory 3.5m  APO APO APO
APO  This is message 447 in the apo35-general archive. You can find
APO  the archive on
APO  To join/leave the list, send mail to
APO  To post a message, mail it to